OpenStudy (anonymous):

solve the quadratic equations? 10x2 – 7x – 12 = 0

OpenStudy (anonymous):

10 (x-7/20)^2-529/40 = 0

OpenStudy (anonymous):

before this steps are (2 x-3) (5 x+4) = 0

OpenStudy (anonymous):

10 x^2 = 7 x+12

OpenStudy (anonymous):

@broderick365 do it

OpenStudy (anonymous):